OpenStudy (anonymous):

Hi I am having trouble with this question.Can anyone help me please? The graph of f (x)₌√x undergoes a vertical stretch by a factor of k, k›1.The vertically stretched graph would look the same if it had been horizontally compressed by a certain factor. Determine this factor

OpenStudy (amistre64):

i think its 1/k

OpenStudy (anonymous):

why can u expalin it?

OpenStudy (amistre64):

maybe i saw that wrong... let me review it

OpenStudy (amistre64):

better examples :) f(16) = ksqrt(16) = 8 ; k = 2 f(16) = sqrt(16k) = 8 16k = 64 k = 4 f(4) = ksqrt(4) = 6 ; k = 3 f(4) = sqrt(4k) = 6 4k = 36 k = 9 these examples suggest that 2k has the same effect as k

OpenStudy (amistre64):

ack .... 2^2 = 4 ; 3^2 = 9 k and k^2 ...

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

youre welcome; I hope you can read thru my lapses in reasoning ;)